I need some help I canot find the correct answer for the following problems
Solve by completing the square 2x^2-8x-11=0
What are the intercepts of y=-x^2+x
Solve by completing the square
x^2-18x+80=0
Is the following trinominal a perfect square
x^2+12+36
Thanks for you assistance

Answer by ankor@dixie-net.com(22740)

Solve by completing the square Solve by completing the square:
2x^2 - 8x - 11 = 0
:
2x^2 - 8x + ___ = + 11
:
Make the coefficient of x^2 = to 1; divide equation by 2
x^2 - 4x + ____ = + 5.5
:
To find the value that completes the square:
Divide the coefficient of x by 2, & square it, (4/2)^2 = 4; add 4 to both sides:
x^2 - 4x + 4 = 5.5 + 4
x^2 - 4x + 4 = 9.5
(x-2)^2 = 9.5
:
Find the square root of both sides:
x - 2 = +/-SqRt(9.5)
:
x = 2 + SqRt(9.5)
and
x = 2 - SqRt(9.5)
:
:
What are the intercepts of:
y = -x^2 + x
-x^2 + x = 0
Factor:
x(-x + 1) = 0
x = 0
and
-x = -1
so
x = +1
The y intercept is 0 because the 3rd term in the quadratic is 0
:
:
Solve by completing the square
x^2 - 18x + 80 = 0
x^2 - 18x + __ = -80
x^2 - 18x + 81 = -80 + 81
(x-9)^2 = +1
x - 9 = +/-SqRt[1]
x = 9 + 1
x = 10
and
x = 9 - 1
x = 8
:
:
Is the following trinominal a perfect square
x^2+12+36
Yes: FOIL (x+6)(x+6)
: